OpenStudy (anonymous):

Please help with limits question. Lim as x approaches 0 , (3/x -3/x^2+x) I know that you use conjugate and the answer is 3, but I can't figure out the steps to get to that answer. Please help!

OpenStudy (solomonzelman):

\(\color{#000000}{\displaystyle \lim_{x\to 0}\left[\frac{3}{x}-\frac{3}{x^2+x}\right]}\) that?

OpenStudy (solomonzelman):

\(\small \color{#000000}{\displaystyle \lim_{x\to 0}\left[\frac{3}{x}-\frac{3}{x^2+x}\right]=\lim_{x\to 0}\left[\frac{3x+3}{x^2+x}-\frac{3}{x^2+x}\right]=\lim_{x\to 0}\left[\frac{3x}{x^2+x}\right]=\lim_{x\to 0}\left[\frac{3}{x+1}\right]}\)

OpenStudy (solomonzelman):

Then, do a direct substitution, plug in x=0 right in!

OpenStudy (anonymous):

thank you thank you thank you!!

OpenStudy (solomonzelman):

\(\color{black}{y(\omega)}\) :)